OpenStudy (anonymous):

Express the complex number in trigonometric form. -6+6 sqrt(3)i

OpenStudy (anonymous):

somebody please help me

OpenStudy (jdoe0001):

gimme one sec, have you found the modulus yet? do you even know what that is?

OpenStudy (anonymous):

the first number i think is 12

OpenStudy (anonymous):

these are the answers

OpenStudy (jdoe0001):

ahemm, one sec

OpenStudy (jdoe0001):

http://mhanswers-auth.mhhe.com/sites/default/files/images/57%20polar.JPG the "r" part is the so-called "modulus", as you can see from the picture is \(\large \sqrt{a^2+b^2}\)

OpenStudy (anonymous):

oh i see

OpenStudy (jdoe0001):

as as you can see from the picture, they way you get the "angle" is by using the tangent function

OpenStudy (anonymous):

well i need help finding the moduous

OpenStudy (jdoe0001):

well, just use => \(\large \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

from the a+bi complex expression

OpenStudy (anonymous):

sqrt((-6)^2+(6 sqrt(3)i)^2 )

OpenStudy (jdoe0001):

right, what does that give you?

OpenStudy (jdoe0001):

$$ r=\sqrt{6^2+(6\sqrt{3})^2} \implies \sqrt{36+36(\sqrt{3})^2} \implies \sqrt{36+36(3)} $$

OpenStudy (jdoe0001):

so, what would that give you?

OpenStudy (anonymous):

i really help guys

OpenStudy (anonymous):

Just use De'Moivre's Theorem.

OpenStudy (anonymous):

ok but how do you find the answer

OpenStudy (anonymous):

these are the answers

OpenStudy (anonymous):

please help me